Write the equation of the line in slope/intercept form.
contains (8,9) and is perpendicular to 2y+9x=-4

Answer by Cromlix(4381)

Put 2y + 9x = -4 into
y = mx + c form.
2y = -9x - 4
y = -9/2x - 2
Gradient/m = -9/2
Using equation:
y - b = m(x - a)
a = 8, b= 9 and m = -9/2
y - 9 = -9/2(x - 8)
y - 9 = -9/2x + 36
y = -9/2x + 36 + 9
y = -9/2x + 45
OR
2y = -9x + 90
Hope this helps.
:-)